Question 3 Using (P)(1-P) P 0<P <1 Evaluate the following integral 4 dy • 1+ Solution
Answer #1

Step 1: Put 1494 - t when yzo ,t=1 when y=0&t=0 dt- - 1 . 443 dy (1+442 1- Ft. csophy a - foto te wys dit + 7, ? (-ve sign asI I & j ci-tt- t ly at Truck: 2 = 1 - 1 1 - 1 I + (1 - 7) - alt Now B[ary] =j ci-t? - 1 t - at I- B [t:1-4 Step 2: B[ary] =.